OpenStudy (anonymous):

Check my work, please. Manda has $1,500 to invest. She invested this money in a certificate of deposit that pays 5.75% over 2 years. How much interest will the certificate deposit pay at the end of the 2-year period? $17.25 $168.75 $172.50 $16.88

OpenStudy (anonymous):

Formula is used to calculate simple interest: I = P * R * T

OpenStudy (sagewilson):

Okay, 5.75 as a decimal would be 0.0575, that multiplied by 2 would be 0.115, then you just multiply that by 1,500.

OpenStudy (anonymous):

I = 1500 * 5.75 % * 2 (500 * 5.75 % * 2 = 57.5)

OpenStudy (anonymous):

So, that'd be a C, then?

OpenStudy (sagewilson):

yeah.

OpenStudy (anonymous):

I appreciate, thank you.

OpenStudy (sagewilson):

no problem.